OpenStudy (anonymous):

I really need help with this. it's a 3 step problem and my lesson teaches absolutely nothing about it. With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in 3 hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. Find the wind speed and the plane's airspeed. {wind speed is ___________a0 mph, airspeed is __________a1 mph}

jimthompson5910 (jim_thompson5910):

let w = windspeed p = speed of plane without any wind at all

jimthompson5910 (jim_thompson5910):

The tailwind speeds the plane up (the tailwind is coming from the tail and it pushes the plane in the direction it wants to go) So with the tailwind, the plane is really going w+p miles per hour

jimthompson5910 (jim_thompson5910):

Distance = Rate * Time d = r*t 1890 = (w+p)*3 that's one equation

jimthompson5910 (jim_thompson5910):

The headwind (wind coming from the head of the plane, pushing it back) slows it down so returning, the headwind makes the plane go p - w mph d = r*t 1890 = (p - w)*3.375 that's the other equation

jimthompson5910 (jim_thompson5910):

now you have a system of two equations in which you can use to solve for p and w